OpenStudy (anonymous):

Got a 70% on a quiz so I want to understand what I got wrong. Click on this question to read the answers what i got wrong.

OpenStudy (anonymous):

Evaluate 2a^3 + (2a)^2 I got 24?

OpenStudy (anonymous):

a = ?

OpenStudy (anonymous):

a=2 sorry

OpenStudy (anonymous):

2^3*2 +4*2^2 = 32

OpenStudy (anonymous):

2(a)^2+4(a^2) 2(2)^2+4(2^2) 8+16=24

OpenStudy (anonymous):

you got 24 wrong

OpenStudy (anonymous):

hahd 2a^3 not 2a^2

myininaya (myininaya):

=32 not 24

OpenStudy (anonymous):

oh my bad blurry computer screen still havent installed software

OpenStudy (anonymous):

Solve for t in the equation B = 9s^2t. a. t=9s^2-B b. t=B-9s^2 c. t=9s^2/B d. t=B/9s^2 I got B?

myininaya (myininaya):

d

OpenStudy (anonymous):

Oh wow I did it 2a^2 not 2a^3

OpenStudy (anonymous):

\[2a^3 + (2a)^2\] \[2\times 2^3+(2\times2)^2\] \[2\times 8+4^2\] \[16+16\] \[32\]

OpenStudy (anonymous):

Solve for c in the equation a=b-c/c

OpenStudy (anonymous):

c=-b/a+1

OpenStudy (anonymous):

c=b/a+1

OpenStudy (anonymous):

c=a+1/b

OpenStudy (anonymous):

hehe stupid mistakes live with them i make the all the time, and face palm myself alot

OpenStudy (anonymous):

c=-a+1/b

OpenStudy (anonymous):

\[a=\frac{b-c}{c}\] \[ca=b-c\] \[ca+c=b\] \[c(a+1)=b\] \[c=\frac{b}{a+1}\]

OpenStudy (anonymous):

c =a+1 --- b

OpenStudy (anonymous):

there seems to be a difference of opinion here. did i write the problem correctly?

OpenStudy (anonymous):

satellite's right sorry got confused with numerator and denominator sorry